A shunt machine is running as a motor off a 500 - V system , taking an armature current of 50 A. if the field current is suddenly increased so as to increase the flux by 20 per cent , calculate the value of the current that would momentarily be fed back into the mains . neglect the shunt current and assume the resistance of the armature circuit to be 0.5 ohms.


1
Expert's answer
2021-03-09T04:54:39-0500

Given quantities:

"U = 500V" "I=50A" "\varPhi_2= 1.2\varPhi_1" "X_L = 0.5\varOmega"

"R = \frac{U}{I} = 10\varOmega" "X_L = wL \to L = \frac{X_L}{w}"

"\varPhi = LI"

"\frac{R}{w}I_2 =1.2* \frac{X_L}{w}I_1 \to I_2= \frac{1.2X_LI_1}{R} = \frac{1.2*0.5*50}{10}=3A"